OpenStudy (anonymous):

A thermometer manufacturer wishes to scale a thermometer in both degrees Celsius and degrees Fahrenheit. Find the missing Celsius degree measures in the illustration. 86°F 59°F 23°F

OpenStudy (anonymous):

C = 5(F-32) / 9 converts F to C for example C = 5(86-32) / 9 = 5* 54 / 9 = 30 degrees C

OpenStudy (anonymous):

ok? just plug in the other values and YOU'RE DONE.

OpenStudy (anonymous):

I now thats the formula but everytime I do the steps I dnt get the answer rite so Im messing up one part

OpenStudy (anonymous):

ok lets do the next one in more detail C = 5(F-32) / 9 substitute F = 59: C = 5(59-32) / 9 always work out whats in the brackets first: 59-32 = 27 s0: C = 5 * 27 / 9 5*27 = = 135 so we have C = 135 / 9 135 divided by 9 = 15 so ts 15 degree centigrade

OpenStudy (anonymous):

I tried to do the last one Im not getting the rite answer I got 5